OpenStudy (anonymous):

Find the distance BF for the points B(1, –1) and F(–3, 5).

OpenStudy (aakashsudhakar):

You know the song, "I Love Rock n' Roll"? Remember these catchy lyrics that go to the tune to that song's chorus... Dis-tance formula! Subtract the X's, Subtract the Y's Dis-tance formula! Square root of the sum of squares! This forms this pretty little equation... D(x, y) = √((x2 - x1)^2 + (y2 - y1)^2) Plug in your values and solve. B(1, -1) --> (x1, y1) F(-3, 5) --> (x2, y2) x1 = 1 x2 = -3 y1 = -1 y2 = 5 D(x, y) = √((-3 - 1)^2 + (5 - -1)^2) D(x, y) = √((-4)^2 + (6)^2) D(x, y) = √(16 + 36) D(x, y) = √(52) D(x, y) = √(13*4) D(x, y) = √(13) * √(4) D(x, y) = 2√(13) The distance between points B and F is equal to exactly 2√(13) units, which is roughly 7.21111111 units of length.